# A Comprehensive Survey for Hyperspectral Image Classification: The Evolution from Conventional to Transformers

### v3

@@ 9 Autoencoders for HSC
The AE is a widely adopted symmetric neural network in HSC for unsupervised feature learning [208, 209, 210]. It prioritizes generating a compressed feature representation of high-dimensional HS data instead of direct classification. The AE architecture typically includes an input layer, a hidden or encoding layer, a reconstruction or decoding layer, and an output layer as shown in Figure 11. During training, the AE learns to encode input data into a latent representation that accurately reconstructs the original input. This is achieved by minimizing the reconstruction error, measuring the difference between the input and output, and enabling effective learning of a compressed feature representation.
@@
The AE is designed to extract essential features from input data, discarding irrelevant information through unsupervised learning. This process yields compact representations, beneficial for tasks like classification in HSC. The Stacked AE (SAE) [209, 211] extends this architecture, stacking multiple AE layers to learn increasingly abstract features. The Denoising AE (DAE) [212], another variant, intentionally corrupts input data with noise, forcing the model to reconstruct the original signal. This enhances the DAE’s ability to handle noisy input, making it useful in scenarios with data corruption. Overall, the AE, SAE, and DAE are powerful tools in HSC for learning compressed, abstract features and handling noisy input.
@@
In Zhu et al.’s [213] approach, multi-layer AEs with maximum noise fraction were combined to learn high-level representations from HS data. Hassanzadeh et al. [214] employed a multi-manifold learning framework combined with the Counteractive AE for unsupervised HSC. Zhang et al. [215] proposed a framework utilizing a Recursive AE (RAE) network to jointly exploit spectral-spatial features, focusing on recursive feature extraction from neighboring pixels based on spectral similarity. Hao et al. [216] introduced a two-stream DNN with a class-specific fusion scheme, incorporating an SAE for spectral features and a CNN for spatial information. Sun et al.’s architecture [217] integrated PCA, guided filters, and sparse AEs with batch-based training to fuse spectral and spatial information. In Zhao et al.’s framework [218], an HSC approach involved selecting spatial resolution from HSIs and utilizing stacked sparse AEs for high-level feature extraction, with classification performed using Random Forest. The hybrid architectures proposed by Sun et al. [217] and Zhao et al. [218] effectively combined PCA, guided filters, and sparse AEs for feature extraction in both spectral and spatial domains.
@@
Wan et al. [219] employed SAEs to extract diverse representations, including spectral-spatial features and multi-fractal features, alongside higher-order statistical representations. Lv et al. [220] proposed a combination of SAEs and ELMs, where features were segmented, transformed using SAEs, and rearranged for input to an ELM-based classifier. Ahmad et al. [55] introduced a computationally efficient multi-layer ELM-based AE for HSC, aiming to enhance feature learning efficiency and effectiveness. Zhou et al. [221] addressed intra-class variability and inter-class similarity challenges in HSC by incorporating a local Fisher discriminant regularization technique for learning compact and discriminative features. Lan et al. [222] combined a k-sparse denoising AE with spectral-restricted spatial features to handle high intra-class variability in spatial features. Paul et al. [223] segmented spectral components in HSI based on mutual information measure, reducing computation time during feature extraction using SAE and incorporating spatial information with Empirical Mode Decomposition (EMD). Liu et al. [224] utilized an SAE-based approach for classifying oil slicks on the sea surface, jointly exploiting spectral-spatial features to improve accuracy.
@@
10 Limitations of Traditional Deep Learning
Applying DL to HSC faces challenges arising from the unique characteristics of HS imaging. These challenges include a large number of narrow spectral bands, high spectral resolution, low spatial resolution, and limited labeled training data. The curse of dimensionality, exacerbated in high-dimensional data scenarios, impacts classification performance, particularly when labeled data is insufficient. The Hughes phenomena, associated with small labeled datasets, can lead to overfitting and poor generalization. The scarcity of labeled HS data, a time-consuming and expensive task, hampers model training. Addressing these challenges requires strategies like feature selection, dimensionality reduction, and augmenting training data through unsupervised or semi-supervised learning. Efficient labeling methods can enhance dataset availability. High intra-class variability, instrumental noise, redundant spectral bands, and spectral mixing further complicate HSC [225]. Researchers are exploring techniques involving domain knowledge, denoising algorithms, feature extraction, and spatial-spectral fusion to tackle these challenges, improving the robustness and accuracy of HSC models. Following are some main challenges that come across when DL is applied to HSC:
Complex Training: Training deep models for HS analysis, belonging to the NP-complete problem class, poses challenges in convergence, especially with a large number of parameters [48, 226]. Optimization techniques like SGD, SGDM, RMSProp, Adam, AdamW, diffGrad, RAdam, gradient centralization, and AngularGrad have addressed this, enhancing training efficiency and convergence for HS analysis [227, 228, 229, 230, 231, 232, 233, 234, 235]. Model Interpretability: The interpretability of DNNs remains challenging due to their black-box nature. Despite efforts to enhance interpretability, DNNs are inherently complex, impacting understanding and decision-making during optimization. Limited Training Data: Supervised deep models in HS imaging face challenges due to limited labeled data, risking overfitting, known as the Hughes phenomena [236, 237]. The high dimensionality of HSI exacerbates this challenge, requiring careful parameter tuning during training [95]. High Computational Burden: DNNs handling large datasets encounter challenges in memory, computation, and storage. Advances in parallel and distributed architectures, along with high-performance computing, enable efficient processing of large datasets in HS imaging [238, 239, 240, 241]. Training Accuracy Degradation: The assumption that deeper networks yield higher accuracy is not always valid. Deeper networks may face challenges like the exploding or vanishing gradient problem, adversely affecting convergence [242, 243]. By mitigating these artifacts and advancing robust techniques, researchers can enhance the reliability and accuracy of HSC, facilitating applications in remote sensing, agriculture, and environmental monitoring.
11 Transformers for HSC
The Transformer architecture, introduced by Vaswani et al. [244] in 2017, has emerged as a powerful DL model for various tasks, such as machine translation, language understanding, text generation, etc. Unlike traditional RNNs or CNNs, the Transformer architecture relies on a self-attention mechanism to capture dependencies between different elements in a sequence, enabling it to effectively model long-range dependencies. This self-attention mechanism allows the model to attend to different parts of the input sequence to build contextualized representations. The Transformer architecture has also found applications in the field of HSC as shown in Figure 12. HSC involves the classification of each pixel in an HSI into different land cover or material classes [12]. The Transformer architecture offers several advantages in this context:
@@ 11 Transformers for HSC
    Capturing Spectral-Spatial Dependencies: The Transformer architecture can effectively capture dependencies between spectral bands and spatial locations in HSIs. By attending to both the spectral and spatial dimensions, the model can learn to extract discriminative features that leverage the rich information present in HS data [245].
@@
    Handling Long-Range Dependencies: HSIs typically exhibit long-range dependencies, where the spectral characteristics of a pixel may depend on distant pixels in both spectral and spatial dimensions [246]. The self-attention mechanism in the Transformer architecture allows the model to capture these long-range dependencies, enabling it to model complex relationships between pixels and make more accurate predictions [181].
@@
    Interpretability: The attention mechanism in the Transformer architecture provides interpretability by highlighting the relevant spectral and spatial features that contribute to the classification decision [247]. This can help in understanding the reasoning process of the model and identifying important features for classification.
@@
In recent studies, researchers have explored different variations of the Transformer architecture for HSC [248, 249, 250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 56, 262, 263, 264, 265, 266, 267, 268, 269, 270, 271, 272, 273], including adaptations like the HS Transformer Network and the Spectral Transformer Network [261]. These models have shown promising results in HSC tasks, achieving competitive performance compared to traditional approaches like CNNs and RNNs. Overall, the Transformer architecture offers a powerful and flexible framework for HSC. Its ability to capture spectral-spatial dependencies, handle long-range dependencies, and generate contextualized representations makes it a promising approach for extracting discriminative features from HS data and improving classification accuracy.
@@
Qing et al. [274] introduced SAT Net, an enhanced Transformer network by incorporating self-attention and spectral attention mechanisms. SAT Net adeptly captures spectral-spatial characteristics. It addresses issues like gradient dispersion and overfitting through the utilization of multi-head self-attention modules and a residual network structure. He et al. [275] proposed a spatial-spectral Transformer (SST) framework. SST integrates a spatial feature extraction CNN with a modified Transformer model known as DenseTransformer to capture sequential correlations within spectral data.
@@
Scheibenreif et al. [276] explored self-supervised masked image reconstruction to enhance Transformer models. Through modifications to the vision Transformer architecture, such as block-wise patch embeddings, spatial-spectral self-attention, spectral positional embeddings, and masked self-supervised pre-training, the study achieves notable accuracy improvements. Additionally, it demonstrates the effectiveness of self-supervised pre-training in enhancing Transformer accuracy, especially in scenarios with limited labeled training data. Qin et al. [277] presented a Transformer-based framework that leverages spatial-spectral-associative contrastive learning to extract spatial and spectral features. This framework employs spatial and spectral Transformer modules to learn high-level semantic features and generates label information through spatial-spectral augmentation and image entropy. Associative optimization integrates cross-domain features, while a Transformer-based classifier addresses class imbalance by incorporating class intersection over union in the loss function.
@@
Ma et al. [278] presented the NEHT network, aiming to overcome the limitations of CNNs in capturing long-range spectral dependencies. The NEHT network integrates a 2D convolution module for dimensionality reduction, stacked CNNs to preserve spectral distribution, and a Transformer self-attention mechanism to capture long-term relationships. It extracts group-wise spatial-spectral characteristics, employs a feature fusion mechanism, and demonstrates superior performance compared to existing methods. Dang et al. [279] proposed the Double-Branch Feature Fusion Transformer (DBFFT) model. This model combines CNNs and Transformers to extract spectral and spatial features, utilizing two attention modules to dynamically adjust the relevance of spectral bands and pixels for classification.
@@
Tan et al. [280] presented a novel method by integrating DL, endmember extraction and purification (EMP), and Transformer models. Unlike CNNs, this approach harnesses the Transformer models’ capacity to capture long-range dependencies and global context information. Initially, EMP features are extracted from the HSI to maximize spatial and spectral information while reducing the band count. Subsequently, a deep network model incorporating Transformer-iN-Transformer (TNT) modules is constructed for end-to-end classification. These modules extract patch- and pixel-level data from the input EMP cubes, effectively utilizing both global and local information.
@@
Hong et al. [261] proposed a Transformer-based backbone network designed for HSC. It includes two main modules: GroupWise spectral embedding (GSE) to capture local spectral differences and cross-layer adaptive fusion (CAF) for enhancing information flow between layers. Similarly, SSFTT model [281] is a DL architecture that integrates a hierarchical CNN module for extracting shallow spatial-spectral features, a Gaussian-weighted feature tokenizer for converting these features into tokenized semantic features, and a Transformer encoder (TE) module for learning deep semantic features.
He et al. [282] proposed Spa-Spe-TR, a two-branch pure Transformer model comprising a spatial Transformer (Spa-TR) for spatial feature extraction and a spectral Transformer (Spe-TR) for capturing spectral dependencies. Their branch fusion technique efficiently combines joint spatial-spectral features, enhancing their relevance and discriminability. Mei et al. [283] introduced GAHT, a Group-Aware Hierarchical Transformer that combines a grouped pixel embedding (GPE) module with a hierarchical Transformer structure to address previous limitations. The GPE module focuses on local interactions within HSI spectral channels, facilitating a global-local approach to HSI categorization. The hierarchical structure gradually reduces the channel size of feature maps to efficiently extract multiscale features, reducing computational costs while maintaining competitive classification results.
@@
The HSI Transformer (HiT) [284] achieved higher accuracy using two modules: the Spectral-Adaptive 3D Convolution Projection (SACP) module and the convolutional Permutator module. The SACP module employs spectral-adaptive 3D convolutional layers to extract both local spatial and long-term spectral information, while the convolutional Permutator module encodes representations in height, width, and spectral dimensions through depthwise and pointwise convolutions. Roy et al. [285] introduced the Spectral-Spatial Morphological Attention Transformer (MorphFormer) model, leveraging learnable spectral and spatial morphological convolutions. MorphFormer captures intrinsic shape information. It incorporates a novel morphological attention mechanism to fuse the CLS token with HSI patch tokens, enriching the abstract representation. Yu et al. [286] introduced MSTNet, a lightweight Transformer-based network that employs an image-based classification architecture to efficiently capture detailed global information. It is adaptable to input images of any size and leverages multi-layer features to enhance the representation of HSIs. Tu et al. [287] proposed LSFAT, a Local Semantic Feature Aggregation-based Transformer architecture. LSFAT enhances the ability of Transformers to represent long-range dependencies of multi-scale features by introducing homogeneous region concepts and neighborhood-aggregation-based embedding (NAE) and attention (NAA) modules. These modules adeptly form multi-scale features and hierarchically capture local spatial semantics. Additionally, LSFAT incorporates a reusable classification token in attention calculation, and a fully connected layer is employed in the final stage for classification after Transformer encoding.
@@
Tang et al. [288] proposed DATE, a Double-Attention Transformer Encoder architecture that efficiently integrates both local spatial and global spectral data by employing two self-attention modules: spectral attention (SPE) and spatial attention (SPA). SPE captures global spectral information from all spectral tokens, while SPA combines local spatial and global spectral properties for each band, facilitating the effective fusion of spectral and spatial semantic data. Zhong et al. [289] introduced SSTN, a Spectral-Spatial Transformer Network, alongside the Factorized Architecture Search (FAS) framework. SSTN incorporates novel spatial attention and spectrum association modules to capture long-range interactions and spectral-spatial correlations in the data efficiently. The FAS framework aids in discovering optimal network designs by decomposing the search process into two distinct subprocedures. This study highlights the advantages of Transformer-based architectures and the effectiveness of factorized architecture search. Zhao et al. [290] proposed a novel framework by combining a multi-attention Transformer (MAT) with an adaptive superpixel segmentation-based AL (ASSAL) technique. The MAT model leverages self-attention and outlook-attention methods to extract both long-range and local contextual variables from input data. To address the challenge of limited labeled samples, the ASSAL technique utilizes an adaptive superpixel segmentation algorithm to identify informative samples for AL. This approach significantly reduces the number of labeled samples required for training while maintaining high classification accuracy.
@@
Moreover, in general, the Swin Transformer (ST) showcases significant strengths, primarily deriving from its hierarchical attention mechanism, which enables the model to efficiently capture information across different scales, analyzing both local and global features [291, 292, 293]. Furthermore, the utilization of windowing-based processing enhances the scalability of STs in contrast to traditional Transformers, facilitating the effective handling of large images with reduced computational complexity [294]. The adaptability of the architecture to diverse tasks makes STs suitable for various applications [295]. In the domain of HSC, STs have demonstrated state-of-the-art (SOTA) performance, outperforming traditional Transformers and CNNs in specific scenarios, underscoring their efficacy for HSC tasks [296, 297]. However, despite these achievements, STs have limitations. While excelling in capturing spatial relationships, they may encounter challenges in dealing with sequential data, rendering tasks reliant on spectral dependencies less optimal for them [298]. Additionally, the hierarchical attention mechanism introduces added complexity. Training large ST models necessitates substantial computational resources, posing challenges for researchers with limited access to high-performance computing. Moreover, akin to other deep models, the interpretability of STs raises concerns, particularly in complex tasks like HSIC, necessitating ongoing research to understand their decision-making processes [293].
@@
Likewise, the vision and spatial-spectral Transformers (SSTs) stand out in their ability to capture global contextual information [299, 300, 301, 302, 303, 304, 305, 306]. The self-attention mechanism enables the model to consider relationships between all HSI regions simultaneously, providing a comprehensive understanding of the visual context [307]. Unlike CNNs, SSTs demonstrate robust scalability to high-resolution HISs, effectively processing large HSI datasets without the need for complex pooling operations. This versatility has led to successful applications of SSTs in HSC, and their adaptable architecture contributes to their broad applicability. Moreover, SSTs reduce reliance on handcrafted features by directly learning hierarchical representations from raw pixel values [261], simplifying the model-building process and often yielding improved performance. Additionally, the attention maps generated by SSTs offer insights into the model’s decision-making process, enhancing interpretability by highlighting the image regions most influential in specific predictions [308].
@@
Despite the accomplishments of SSTs, they do come with certain limitations. Notably, training large SSTs can pose significant computational challenges, particularly as the model size increases [246, 91]. The self-attention mechanism introduces quadratic complexity with respect to sequence length, potentially impeding scalability [309]. Unlike CNNs, which inherently possess translation invariance through shared weight convolutional filters, SSTs may struggle to capture spatial relationships that remain invariant to small translations in the input [310]. Moreover, SSTs rely on dividing input images into fixed-size patches during the tokenization process, which may not efficiently capture fine-grained details [311, 312]. The quadratic scaling of self-attention poses challenges, particularly when dealing with long sequences. Additionally, achieving optimal performance with SSTs often requires substantial amounts of training data. Attempting to train these models on smaller datasets may lead to overfitting, thus limiting their effectiveness in scenarios with limited labeled data [12].

### v4

@@ 9 Autoencoders for HSC
The AE is a widely adopted symmetric neural network in HSC for unsupervised feature learning (Tian et al., 2024a, ; Li et al., 2024g, ; Huo et al.,, 2024). It prioritizes generating a compressed feature representation of high-dimensional HS data instead of direct classification. The AE architecture typically includes an input layer, a hidden or encoding layer, a reconstruction or decoding layer, and an output layer as shown in Figure 11. During training, the AE learns to encode input data into a latent representation that accurately reconstructs the original input. This is achieved by minimizing the reconstruction error, measuring the difference between the input and output, and enabling effective learning of a compressed feature representation.
@@
The AE is designed to extract essential features from input data, discarding irrelevant information through unsupervised learning. This process yields compact representations, beneficial for tasks like classification in HSC. The Stacked AE (SAE) (Cao et al., 2024a, ) extends this architecture, stacking multiple AE layers to learn increasingly abstract features. The Denoising AE (DAE) (Zhao et al., 2024c, ), another variant, intentionally corrupts input data with noise, forcing the model to reconstruct the original signal. This enhances the DAE’s ability to handle noisy input, making it useful in scenarios with data corruption. Overall, the AE, SAE, and DAE are powerful tools in HSC for learning compressed, abstract features and handling noisy input.
@@
In Zhu et al.’s (Zhu et al.,, 2017) approach, multi-layer AEs with maximum noise fraction were combined to learn high-level representations from HS data. Hassanzadeh et al. (Hassanzadeh et al.,, 2017) employed a multi-manifold learning framework combined with the Counteractive AE for unsupervised HSC. Zhang et al. (Zhang et al., 2017b, ) proposed a framework utilizing a Recursive AE (RAE) network to jointly exploit spectral-spatial features, focusing on recursive feature extraction from neighboring pixels based on spectral similarity. Hao et al. (Hao et al.,, 2017) introduced a two-stream DNN with a class-specific fusion scheme, incorporating an SAE for spectral features and a CNN for spatial information. Sun et al.’s architecture (Sun et al.,, 2017) integrated PCA, guided filters, and sparse AEs with batch-based training to fuse spectral and spatial information. In Zhao et al.’s framework (Zhao et al.,, 2017), an HSC approach involved selecting spatial resolution from HSIs and utilizing stacked sparse AEs for high-level feature extraction, with classification performed using Random Forest. The hybrid architectures proposed by Sun et al. (Sun et al.,, 2017) and Zhao et al. (Zhao et al.,, 2017) effectively combined PCA, guided filters, and sparse AEs for feature extraction in both spectral and spatial domains.
@@
Wan et al. (Wan et al.,, 2017) employed SAEs to extract diverse representations, including spectral-spatial features and multi-fractal features, alongside higher-order statistical representations. Lv et al. (Lv et al.,, 2017) proposed a combination of SAEs and ELMs, where features were segmented, transformed using SAEs, and rearranged for input to an ELM-based classifier. Ahmad et al. (Ahmad et al., 2019c, ) introduced a computationally efficient multi-layer ELM-based AE for HSC, aiming to enhance feature learning efficiency and effectiveness. Zhou et al. (Zhou et al.,, 2019) addressed intra-class variability and inter-class similarity challenges in HSC by incorporating a local Fisher discriminant regularization technique for learning compact and discriminative features. Lan et al. (Lan et al.,, 2019) combined a k-sparse denoising AE with spectral-restricted spatial features to handle high intra-class variability in spatial features. Paul et al. (Paul and Kumar,, 2018) segmented spectral components in HSI based on mutual information measure, reducing computation time during feature extraction using SAE and incorporating spatial information with Empirical Mode Decomposition (EMD). Liu et al. (Liu et al., 2019a, ) utilized an SAE-based approach for classifying oil slicks on the sea surface, jointly exploiting spectral-spatial features to improve accuracy.
@@
10 Limitations of Traditional Deep Learning
Applying DL to HSC faces challenges arising from the unique characteristics of HS imaging. These challenges include a large number of narrow spectral bands, high spectral resolution, low spatial resolution, and limited labeled training data. The curse of dimensionality, exacerbated in high-dimensional data scenarios, impacts classification performance, particularly when labeled data is insufficient. The Hughes phenomena, associated with small labeled datasets, can lead to overfitting and poor generalization. The scarcity of labeled HS data, a time-consuming and expensive task, hampers model training. Addressing these challenges requires strategies like feature selection, dimensionality reduction, and augmenting training data through unsupervised or semi-supervised learning. Efficient labeling methods can enhance dataset availability. High intra-class variability, instrumental noise, redundant spectral bands, and spectral mixing further complicate HSC (Hong et al.,, 2018). Researchers are exploring techniques involving domain knowledge, denoising algorithms, feature extraction, and spatial-spectral fusion to tackle these challenges, improving the robustness and accuracy of HSC models. Following are some main challenges that come across when DL is applied to HSC:
Complex Training: Training deep models for HS analysis, belonging to the NP-complete problem class, poses challenges in convergence, especially with a large number of parameters (Ahmad et al., 2021b, ; Nguyen and Hein,, 2017). Optimization techniques like SGD, SGDM, RMSProp, Adam, AdamW, diffGrad, RAdam, gradient centralization, and AngularGrad have addressed this, enhancing training efficiency and convergence for HS analysis (Dubey et al.,, 2019; Liu et al., 2019c, ; Yong et al.,, 2020). Model Interpretability: The interpretability of DNNs remains challenging due to their black-box nature. Despite efforts to enhance interpretability, DNNs are inherently complex, impacting understanding and decision-making during optimization. Limited Training Data: Supervised deep models in HS imaging face challenges due to limited labeled data, risking overfitting, known as the Hughes phenomena (Erhan et al.,, 2010; Dutta et al.,, 2013). The high dimensionality of HSI exacerbates this challenge, requiring careful parameter tuning during training (Paoletti et al.,, 2019). High Computational Burden: DNNs handling large datasets encounter challenges in memory, computation, and storage. Advances in parallel and distributed architectures, along with high-performance computing, enable efficient processing of large datasets (Alom et al.,, 2019). Training Accuracy Degradation: The assumption that deeper networks yield higher accuracy is not always valid. Deeper networks may face challenges like the exploding or vanishing gradient problem, adversely affecting convergence (He et al.,, 2016; Bengio et al.,, 1994). By mitigating these artifacts and advancing robust techniques, researchers can enhance the reliability and accuracy of HSC, facilitating applications in remote sensing, agriculture, and environmental monitoring.
11 Transformers for HSC
The Transformer architecture, introduced by Vaswani et al. (Vaswani et al.,, 2017) in 2017, has emerged as a powerful DL model for various tasks, such as machine translation, language understanding, text generation, etc. Unlike traditional RNNs or CNNs, the Transformer architecture relies on a self-attention mechanism to capture dependencies between different elements in a sequence, enabling it to effectively model long-range dependencies. This self-attention mechanism allows the model to attend to different parts of the input sequence to build contextualized representations. The Transformer architecture has also found applications in the field of HSC as shown in Figure 12. HSC involves the classification of each pixel in an HSI into different land cover or material classes (Ahmad et al., 2024d, ). The Transformer architecture offers several advantages in this context:
@@ 11 Transformers for HSC
    Capturing Spectral-Spatial Dependencies: The Transformer architecture can effectively capture dependencies between spectral bands and spatial locations in HSIs. By attending to both the spectral and spatial dimensions, the model can learn to extract discriminative features that leverage the rich information present in HS data (Chen et al., 2024b, ).
@@
    Handling Long-Range Dependencies: HSIs typically exhibit long-range dependencies, where the spectral characteristics of a pixel may depend on distant pixels in both spectral and spatial dimensions Fang et al., (2024). The self-attention mechanism in the Transformer architecture allows the model to capture these long-range dependencies, enabling it to model complex relationships between pixels and make more accurate predictions (Zhao et al., 2024b, ).
@@
    Interpretability: The attention mechanism in the Transformer architecture provides interpretability by highlighting the relevant spectral and spatial features that contribute to the classification decision Böhle et al., (2024). This can help in understanding the reasoning process of the model and identifying important features for classification.
@@
In recent studies, researchers have explored different variations of the Transformer architecture for HSC (Ahmad et al., 2024a, ; Butt et al.,, 2024; Ghosh et al.,, 2022; Li et al., 2024e, ; Hong et al., 2020d, ), including adaptations like the HS Transformer Network and the Spectral Transformer Network. Transformer models have shown promising results in HSC tasks, achieving competitive performance compared to traditional approaches like CNNs and RNNs (Yu et al., 2024b, ; He and Chen,, 2019; Bai et al.,, 2022). Overall, the Transformer architecture offers a powerful and flexible framework for HSC (Ahmad et al., 2024e, ). Its ability to capture spectral-spatial dependencies, handle long-range dependencies, and generate contextualized representations makes it a promising approach for extracting discriminative features from HS data and improving classification accuracy.
@@
Qing et al. (Qing et al.,, 2021) introduced SAT Net, an enhanced Transformer network by incorporating self-attention and spectral attention mechanisms. SAT Net adeptly captures spectral-spatial characteristics. It addresses issues like gradient dispersion and overfitting through the utilization of multi-head self-attention modules and a residual network structure. He et al. (He et al.,, 2021) proposed a spatial-spectral Transformer (SST) framework. SST integrates a spatial feature extraction CNN with a modified Transformer model known as DenseTransformer to capture sequential correlations within spectral data.
@@
Scheibenreif et al. (Scheibenreif et al.,, 2023) explored self-supervised masked image reconstruction to enhance Transformer models. Through modifications to the vision Transformer architecture, such as block-wise patch embeddings, spatial-spectral self-attention, spectral positional embeddings, and masked self-supervised pre-training, the study achieves notable accuracy improvements. Additionally, it demonstrates the effectiveness of self-supervised pre-training in enhancing Transformer accuracy, especially in scenarios with limited labeled training data. Qin et al. (Qin and Zhao,, 2023) presented a Transformer-based framework that leverages spatial-spectral-associative contrastive learning to extract spatial and spectral features. This framework employs spatial and spectral Transformer modules to learn high-level semantic features and generates label information through spatial-spectral augmentation and image entropy. Associative optimization integrates cross-domain features, while a Transformer-based classifier addresses class imbalance by incorporating class intersection over union in the loss function.
@@
Ma et al. (Ma et al.,, 2022) presented the NEHT network, aiming to overcome the limitations of CNNs in capturing long-range spectral dependencies. The NEHT network integrates a 2D convolution module for dimensionality reduction, stacked CNNs to preserve spectral distribution, and a Transformer self-attention mechanism to capture long-term relationships. It extracts group-wise spatial-spectral characteristics, employs a feature fusion mechanism, and demonstrates superior performance compared to existing methods. Dang et al. (Dang et al.,, 2023) proposed the Double-Branch Feature Fusion Transformer (DBFFT) model. This model combines CNNs and Transformers to extract spectral and spatial features, utilizing two attention modules to dynamically adjust the relevance of spectral bands and pixels for classification.
@@
Tan et al. (Tan et al.,, 2021) presented a novel method by integrating DL, endmember extraction and purification (EMP), and Transformer models. Unlike CNNs, this approach harnesses the Transformer models’ capacity to capture long-range dependencies and global context information. Initially, EMP features are extracted from the HSI to maximize spatial and spectral information while reducing the band count. Subsequently, a deep network model incorporating transformer-iN-transformer (TNT) modules is constructed for end-to-end classification. These modules extract patch- and pixel-level data from the input EMP cubes, effectively utilizing both global and local information.
@@
Hong et al. (Hong et al.,, 2022) proposed a Transformer-based backbone network designed for HSC. It includes two main modules: GroupWise spectral embedding (GSE) to capture local spectral differences and cross-layer adaptive fusion (CAF) for enhancing information flow between layers. He et al. (He et al.,, 2022) proposed Spa-Spe-TR, a two-branch pure Transformer model comprising a spatial Transformer (Spa-TR) for spatial feature extraction and a spectral Transformer (Spe-TR) for capturing spectral dependencies. Their branch fusion technique efficiently combines joint spatial-spectral features, enhancing their relevance and discriminability.
@@
The HSI Transformer (HiT) (Yang et al.,, 2022) achieved higher accuracy using two modules: the Spectral-Adaptive 3D Convolution Projection (SACP) module and the convolutional Permutator module. The SACP module employs spectral-adaptive 3D convolutional layers to extract both local spatial and long-term spectral information, while the convolutional Permutator module encodes representations in height, width, and spectral dimensions through depthwise and pointwise convolutions. Roy et al. (Roy et al.,, 2023) introduced the Spectral-Spatial Morphological Attention Transformer (MorphFormer) model, leveraging learnable spectral and spatial morphological convolutions. MorphFormer captures intrinsic shape information. It incorporates a novel morphological attention mechanism to fuse the CLS token with HSI patch tokens, enriching the abstract representation. Tu et al. (Tu et al.,, 2022) proposed LSFAT, a Local Semantic Feature Aggregation-based Transformer architecture. LSFAT enhances the ability of Transformers to represent long-range dependencies of multi-scale features by introducing homogeneous region concepts and neighborhood-aggregation-based embedding (NAE) and attention (NAA) modules. These modules adeptly form multi-scale features and hierarchically capture local spatial semantics. Additionally, LSFAT incorporates a reusable classification token in attention calculation, and a fully connected layer is employed in the final stage for classification after Transformer encoding.
@@
Tang et al. (Tang et al.,, 2023) proposed DATE, a Double-Attention Transformer Encoder architecture that efficiently integrates both local spatial and global spectral data by employing two self-attention modules: spectral attention (SPE) and spatial attention (SPA). SPE captures global spectral information from all spectral tokens, while SPA combines local spatial and global spectral properties for each band, facilitating the effective fusion of spectral and spatial semantic data. Zhong et al. (Zhong et al., 2021b, ) introduced SSTN, a Spectral-Spatial Transformer Network, alongside the Factorized Architecture Search (FAS) framework. SSTN incorporates novel spatial attention and spectrum association modules to capture long-range interactions and spectral-spatial correlations in the data efficiently. The FAS framework aids in discovering optimal network designs by decomposing the search process into two distinct subprocedures. This study highlights the advantages of Transformer-based architectures and the effectiveness of factorized architecture search. Zhao et al. (Zhao et al.,, 2023) proposed a novel framework by combining a multi-attention Transformer (MAT) with an adaptive superpixel segmentation-based AL (ASSAL) technique. The MAT model leverages self-attention and outlook-attention methods to extract both long-range and local contextual variables from input data. To address the challenge of limited labeled samples, the ASSAL technique utilizes an adaptive superpixel segmentation algorithm to identify informative samples for AL. This approach significantly reduces the number of labeled samples required for training while maintaining high classification accuracy.
@@
Moreover, in general, the Swin Transformer (ST) showcases significant strengths, primarily deriving from its hierarchical attention mechanism, which enables the model to efficiently capture information across different scales, analyzing both local and global features (Huang et al.,, 2022; Zu et al.,, 2024; Liu et al., 2023a, ). Furthermore, the utilization of windowing-based processing enhances the scalability of STs in contrast to traditional Transformers, facilitating the effective handling of large images with reduced computational complexity (Farooque et al.,, 2023). The adaptability of the architecture to diverse tasks makes STs suitable for various applications (Xie et al.,, 2023). In the domain of HSC, STs have demonstrated state-of-the-art (SOTA) performance, outperforming traditional Transformers and CNNs in specific scenarios, underscoring their efficacy for HSC tasks (Ayas and Tunc-Gormus,, 2022; Peng et al.,, 2023). However, despite these achievements, STs have limitations. While excelling in capturing spatial relationships, they may encounter challenges in dealing with sequential data, rendering tasks reliant on spectral dependencies less optimal for them (Long et al.,, 2023). Additionally, the hierarchical attention mechanism introduces added complexity. Training large ST models necessitates substantial computational resources, posing challenges for researchers with limited access to high-performance computing. Moreover, akin to other deep models, the interpretability of STs raises concerns, particularly in complex tasks like HSIC, necessitating ongoing research to understand their decision-making processes (Liu et al., 2023a, ).
@@
Likewise, the vision and spatial-spectral Transformers (SSTs) stand out in their ability to capture global contextual information (Wang et al., 2024c, ; Wang et al., 2024b, ; Hou et al.,, 2024). The self-attention mechanism enables the model to consider relationships between all HSI regions simultaneously, providing a comprehensive understanding of the visual context (Mei et al.,, 2022). Unlike CNNs, SSTs demonstrate robust scalability to high-resolution HISs, effectively processing large HSI datasets without the need for complex pooling operations. This versatility has led to successful applications of SSTs in HSC, and their adaptable architecture contributes to their broad applicability. Moreover, SSTs reduce reliance on handcrafted features by directly learning hierarchical representations from raw pixel values (Hong et al.,, 2022), simplifying the model-building process and often yielding improved performance. Additionally, the attention maps generated by SSTs offer insights into the model’s decision-making process, enhancing interpretability by highlighting the image regions most influential in specific predictions (Chen et al., 2024a, ).
@@
Despite the advancements, SSTs do come with certain limitations. Notably, training large SSTs can pose significant computational challenges, particularly as the model size increases (Fang et al.,, 2024; Ye et al.,, 2024). The self-attention mechanism introduces quadratic complexity with respect to sequence length, potentially impeding scalability (Huang et al.,, 2024). Unlike CNNs, which inherently possess translation invariance through shared weight convolutional filters, SSTs may struggle to capture spatial relationships that remain invariant to small translations in the input (Sun et al.,, 2024). Moreover, SSTs rely on dividing input images into fixed-size patches during the tokenization process, which may not efficiently capture fine-grained details (Kim et al.,, 2024; Shi et al.,, 2024). The quadratic scaling of self-attention poses challenges, particularly when dealing with long sequences. Additionally, achieving optimal performance with SSTs often requires substantial amounts of training data. Attempting to train these models on smaller datasets may lead to overfitting, thus limiting their effectiveness in scenarios with limited labeled data (Ahmad et al., 2024d, ).